OpenStudy (anonymous):

hi can someone help me i will fan you and give you a medal anyways Factor this expression completely, then place the factors in the proper location on the grid. 1 - 4y2 i got -(2y+1(2y-1) but it does not say its correct\ thank you bye

OpenStudy (anonymous):

thanks

OpenStudy (pokemon23):

No problem

OpenStudy (mathmale):

@pokemon23: I beg your pardon, but we DO have a problem here: You are doing someone else's work and providing free answers. Surely you know by know, if you've read the Code of Conduct, that you are expected to help others find their OWN answers, not to provide them.

OpenStudy (pokemon23):

My fault I haven't log in for a while.

OpenStudy (mathmale):

All right. Thank you for recognizing what I've said. Challenge yourself to find creative ways of helping others find their own answers without actually providing those answers yourself.

OpenStudy (mathmale):

@teamakesmepee : Look at \[1 - 4y^2\]as the difference of two squares. Note that 1=(1)^2 and that (2y)^2. Also note that there's a "special product," a^2 - b^2, which easily factors into (a-b)(a+b). Please use this info to factor \[1 - 4y^2 = (1)^2 - (2y)^2\]

OpenStudy (anonymous):

thank you @mathmale